SubjectCover

Question 79 of 106 in Unit 8

Handling Data 5

3 marks · Easy

The table shows the probability that a spinner will land on each colour.

ColourProbability
Red0.3
Green0.5
Bluep

a) Write down the value of .

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

0.2 (±0.01)

Common mistakes: Adding instead of subtracting; Writing 1.0 or -0.2

Show mark scheme and worked solution

Marking points

M1 · 1 mark · States or uses the rule that the sum of all probabilities is 1.

M2 · 1 mark · Substitutes known probabilities correctly into the rule: 0.3 + 0.5 + p = 1 or 0.3 + 0.5 = 0.8.

A1 · 1 mark · Finds correct value: p = 0.2.

Worked solution

The probability of all possible spinner outcomes must add up to 1.

Add the given probabilities together:

So, p = 1 - 0.8 = 0.2

Report a problem with this question
What's wrong?